# Live Music Models

### v2

@@ 1 Introduction
Music exists in two complementary forms: as static recorded pieces (“music as a noun”), and as live performances collectively experienced in real time (“music as a verb”) Small (1998). This second form of live music is particularly tied to the fundamental human experiences of creative flow Wrigley and Emmerson (2013); Chirico et al. (2015), embodied expression Jensenius (2022), and social connection Trost et al. (2024). Despite this, modern generative AI systems for musical audio have had an overwhelming emphasis on offline, turn-based generation Dhariwal et al. (2020); Forsgren and Martiros (2022); Agostinelli et al. (2023); Copet et al. (2023); Donahue et al. (2023); Chen et al. (2024); Evans et al. (2024a, b, 2025); Yuan et al. (2025).
Live music represents a new frontier for generative AI, one with numerous opportunities and technical challenges. In the conventional offline setting, users input control information, wait $L$ seconds (offline latency), and receive $T$ seconds of audio. In our proposed live setting, users continuously input control information, receiving $T$ seconds of an uninterrupted audio stream from $T$ seconds of interaction, with $D$ seconds of delay between their control inputs and their influence on the audio stream. Placing users in a continuous perception-action loop promotes more active creation, creates higher-bandwidth interaction, fosters personalized expression and emphasize the process as much as the product. However, meeting these rigid synchronization requirements while maintaining high quality audio generation is a challenging task for machine learning models.
@@
Open-weights models are particularly well-suited for live generative music because they can run locally on users’ devices. On-device inference has numerous benefits in music (Zhou et al., 2024), enabling (1) lower latency by eliminating network requests, (2) higher reliability, facilitating usage in real-world contexts, (3) privacy guarantees, and (4) customization by artists. Cloud-based APIs address complementary use cases, offering access to models running on specialized hardware that are more powerful than those that would run on edge devices, at the cost of some of the benefits of open-weights models.
To allow users to navigate these tradeoffs based on their application goals, we introduce a pair of systems that span both paradigms: Magenta RT (open-weights, on-device) and Lyria RT (API, cloud-based). Both use the same core methodological framework, which centers around codec language modeling (Van Den Oord et al., 2017; Wu et al., 2024) (Figure 1). Specifically, we train a language model (LM) to generate audio tokens from SpectroStream (Li et al., 2025), using a method similar to (Yang et al., 2023; Défossez et al., 2024) to achieve live streaming (McWilliams et al., 2025).
We focus our subsequent discussion primarily on Magenta RT, which may be of higher interest to the AI research community for its ability to be finetuned with new controls (Lin et al., 2023) or explored for transfer learning (Castellon et al., 2021). Magenta RT offers first-of-its-kind live generation among open-weights models. On automatic metrics of music quality, it outperforms existing offline open-weights music generation models like MusicGen (Large) (Copet et al., 2023) and Stable Audio Open Evans et al. (2025). Moreover, Magenta RT uses 3838% fewer parameters (750750M) than Stable Audio Open (1.21.2B), and 7777% fewer than MusicGen Large (3.33.3B). Along with our codebase and model weights, we release a set of demos that run in real time on free-tier Colab TPUs (v2-8) and showcase three distinct use cases: live generation, finetuning, and a novel live audio input interaction that we call audio injection (Section 4.2).  [^3] Audio samples and links to code, weights and demos are provided in the online supplement:
@@
Figure 1: Magenta RealTime is a live music model that generates an uninterrupted stream of music and responds continuously to user input. It generates audio in two-second chunks using a pipeline with three components: (1) MusicCoCa, a style embedding model, (2) SpectroStream (Li et al., 2025), an audio codec model, and (3) an encoder-decoder language model. For each chunk, a style embedding is computed via a weighted average of MusicCoCa embeddings of text and audio prompts from the user. Given this style embedding and $10$ seconds ($5$ chunks) of past audio context, the language model decoder generates SpectroStream audio tokens for the new chunk, which is then decoded to audio.
@@ 2 Method
Magenta RT is a codec language model (Van Den Oord et al., 2017; Wu et al., 2024) designed to generate high-fidelity stereo audio in real time based on acoustic style conditioning provided by a user. To achieve this, we adopt a pipeline-based approach, using an encoder-decoder Transformer (Vaswani et al., 2017) to model audio tokens from SpectroStream (Li et al., 2025) conditioned on style embeddings from our proposed MusicCoCa (Figure 1). See Appendix A for related work.
@@ 2.1 Audio tokenization via SpectroStream
Here we adopt the recently proposed SpectroStream codec Li et al. (2025), a full-band ($f_{s}=48$kHz) multi-channel neural audio codec based on residual vector quantization (RVQ) (Zeghidour et al., 2021), with an overall bandwidth of 1616kbps ($f_{k}=25$Hz, $d_{c}=64$, $|\mathbb{V}_{c}|=1024$). To facilitate live streaming, we reduce the bandwidth to 44kbps by generating only the first $16$ RVQ levels (coarse and medium from Figure 3), yielding a live throughput target of $400$ tokens per second. See Section C.1 for more details.
@@ 2.2 Style embeddings via MusicCoCa
Our embedding model, MusicCoCa, builds upon MuLan Huang et al. (2022) and CoCa Yu et al. (2022). It is a contrastive captioner (CoCa) consisting of two embedding towers, mapping each modality to a shared $768$-dimensional space. The audio embedding tower $M_{A}$ is a $12$-layer VisionTransformer (ViT) Dosovitskiy et al. (2021). Its input is a log-mel spectrogram of a $10$s slice of 1616kHz audio ($128$ channels and length $992$; split into patches of size $16\times 16$). The text embedding tower $M_{T}$ is a $12$-layer Transformer, which operates on tokenized text with a maximum sequence length of $128$ tokens. We use attention pooling to reduce the activations of each tower to a single 768768d embedding, which can be subsequently quantized into $12$ discrete tokens with codebook size $|\mathbb{V}_{m}|=1024$. This tokenized representation (of audio or text) is then used as a conditioning signal in the LM encoder. Variable-length audio of duration $T$ may be embedded by zero padding to a multiple of $10$ seconds and then mean pooling across $\lceil\frac{T}{10}\rceil$ chunks.
In addition to the two embedding towers, MusicCoCa has a text decoder which can generate audio text captions. In our application this decoder is a shallow 3-layer Transformer which only serves a regularizing purpose. The MusicCoCa optimization objective, based on the CoCa framework Yu et al. (2022), consists of contrastive and generative loss components which we weigh equally. We train the model using the Adafactor optimizer with a learning rate of $1\times 10^{-4}$ and 1,0001{,}000 warmup steps for a total of 16,00016{,}000 steps. The batch size is $1024$.
@@ 2.3 Modeling Framework
Magenta RT operates on audio tokens from a discrete audio codec, following an established practice in audio modeling Van Den Oord et al. (2017); Borsos et al. (2023); Agostinelli et al. (2023); Copet et al. (2023). Our model autoregressively predicts discrete audio tokens conditioned on both preceding audio and a shared audio-text embedding. This modeling mechanism partly mirrors the MusicLM architecture Agostinelli et al. (2023), itself based on AudioLM Borsos et al. (2023). Similarly to MusicLM, we use a Transformer-based architecture Vaswani et al. (2017) for the token prediction model and enable text conditioning by leveraging a joint audio-text embedding model, using the target audio signal at training time, and text inputs at inference time. To facilitate live generation, we propose two high-level changes relative to MusicLM: (1) we replace the hierarchical cascade of multiple LMs with a single LM, using a recent method (McWilliams et al., 2025) similar to (Yang et al., 2023; Défossez et al., 2024) for efficiency, and (2) we propose a chunk-based autogressive approach to allow for infinite streaming generation.
@@ 2.3.1 Chunk-based autoregression with coarse context
In order to generate an infinite stream of audio, at inference time we must be able to predict an audio sequence with a length likely larger than what the model has seen during training. Such a mismatch in sequence length between training and inference is often found to result in unpredictable behavior and degraded performance Vaswani et al. (2017); Huang et al. (2018). Previous work has addressed this issue via sliding attention windows with a relative positional encoding scheme Su et al. (2024); Press et al. (2021), informing the model about the relative distance between tokens instead of describing their absolute position within the sequence.
@@ 2.4 Encoder-Decoder Language Model
To model this distribution, we use an encoder-decoder Transformer Vaswani et al. (2017) LM trained with T5X Raffel et al. (2020); Roberts et al. (2023). We release pre-trained models using the T5  (Raffel et al., 2020) Base and Large configurations.
@@
A key differentiating factor compared to prior work is our imposed constraint of achieving RTF $\geq 1\times$ generation to enable live interactive applications. Past work such as MusicLM (Agostinelli et al., 2023) proposes a hierarchical cascade of language models to model tokens efficiently, while MusicGen (Copet et al., 2023) proposes a delay pattern—neither approach would achieve RTF $\geq 1\times$ for full bandwidth audio tokens. Instead, based on McWilliams et al. (2025), our decoder comprises two connected Transformer modules. The “temporal” module constructs a temporal context by processing acoustic frames, where RVQ tokens within each frame are embedded and aggregated to yield a single frame-level embedding. The “depth” module then performs autoregressive prediction of the RVQ indices conditioned on the previous temporal context. With this setup, we achieve RTF=$1.8$ on H100 GPU with the T5 Large configuration.
@@ 3 Experiments
Magenta RT, and live music models more broadly, enable new types of interaction that are best assessed through direct human evaluation requiring extensive use. As such, many of the design choices in Magenta RT were evaluated through play testing by team members and partner musicians, optimizing for subjective feeling of how expressive and engaging the resulting instrument is to play.
We complement this subjective measure with more constrained experiments to examine the effects of our controls and provide comparisons to existing models where possible. We focus these experiments on assessing core capabilities that are common to both live and offline music audio generation models, such as audio quality and adherence to text conditioning (Section 3.2.1), alongside others that are unique to our live music models, such as the ability to generate musical transitions following changes in the conditioning signal (Section 3.2.2).
@@ 3.1 Experimental Set-up
For inference, prompts (text or audio) are embedded and tokenized by MusicCoca using the first $6$ RVQ levels to match training. We sample with classifier-free guidance (CFG) Ho and Salimans (2021); Kreuk et al. (2023); Sanchez et al. (2023), using a temperature of $1.3$, a top-K of $40$, and a CFG weight of $5.0$.
@@ 3.2 Results
Table 1: Instrumental music generation results on the Song Describer Dataset Manco et al. (2023). We compare to open models, using a fixed length of $47$s for all samples, though our models are capable of arbitrary-length generation. For all prior models, we report results from Evans et al. (2025).
@@
Stable Audio Open Evans et al. (2025) 44.1 kHz 1.1B 96.51 0.55 0.41
MusicGen-stereo-large Copet et al. (2023) 32 kHz 3.3B 190.47 0.52 0.31
@@ 3.2.1 Audio quality and adherence to fixed text prompts
In this section, we compare Magenta RT to prior work under the offline text-to-music generation setting, where we keep the text conditioning fixed and sequentially generate chunks of audio up to a target length of $47$ seconds. While our model can generate audio of arbitrary length, we choose this duration for a fair comparison between models. Following the evaluation protocol in Evans et al. (2025), we then assess the quality and text adherence of the resulting generations using three established metrics, the Fréchet Distance based on OpenL3 embeddings Cramer et al. (2019) (FDopenl3), the Kullback–Leibler divergence (KLpasst) and CLAPscore.
We show the results in Table 1. Magenta RT has the lowest FDopenl3 and KLpasst scores, indicating that the generated audio is plausible and closely matches the eval reference audio, including at the level of semantic correspondence Evans et al. (2025). The CLAPscore measures how well generated audio adheres to the specified text prompt. On this metric, Magenta RT scores between the other two models. The higher score of Stable Audio Open may be related to the fact that their model uses CLAP embeddings during training, as opposed to our model which trains using MusicCoCa.
@@ 3.2.2 Generating musical transitions
Live music models unlock the capability of responding dynamically to user inputs. To test this, we create a prompt transition evaluation. We pick pairs of text prompts and task the model with creating musical transitions between them. We linearly interpolate between MusicCoCa text embeddings of the start and end prompt over $60$ seconds ($6$ steps, $10$ seconds//step), and use this as style conditioning. We then measure cosine similarity between the audio embedding of the outputs and the conditioning embedding at that time. We perform transitions for $128$ prompt pairs (Appendix F).
@@ Appendix A Related Work
Our work on live music models connects to longstanding goals in computer music of live, computer-aided performance. Earlier work centered around score following: tracking live performances against known musical material to enable computer accompaniment (Dannenberg, 1984; Vercoe, 1984; Orio et al., 2003), or systems that could both track input from human musicians and generate new symbolic material using simple probabilistic models (Lewis, 2000; Thom, 2000; Pachet, 2003; Kondak et al., 2016). More recently, numerous systems have proposed live interaction with generative AI models of symbolic music (Donahue et al., 2019; Jiang et al., 2020; Benetatos et al., 2020; Blanchard et al., 2024; Wu et al., 2025a), or small models trained on narrow music audio distributions (Engel et al., 2020; Caillon and Esling, 2021). Here we aim to bridge the gap between the live interaction paradigm and the broad audio generation and control capabilities of large-scale generative AI models of music audio.
Many offline music audio generation models are based off of codec LMs (Van Den Oord et al., 2017; Wu et al., 2024). Codec LMs are theoretically capable of streaming provided they meet two criteria: (1) the language model and codec are causal (outputs never depend on future inputs), and (2) they generate with RTF $\geq 1\times$ for some chunk size $C$ (which lower bounds the control delay $D$). To the best of our knowledge, no existing codec LM music generation model (Dieleman et al., 2018; Dhariwal et al., 2020; Agostinelli et al., 2023; Copet et al., 2023; Yuan et al., 2025; Liu et al., 2025; Gong et al., 2025) satisfies both criteria (most use non-causal codecs). Other music generation systems (Forsgren and Martiros, 2022; Chen et al., 2024; Evans et al., 2024a, 2025, b) are based off of latent diffusion—many achieve a throughput RTF $\geq 1\times$ but generate in a non-causal fashion. Some “outpainting” methods have been proposed for latent diffusion music models (Levy et al., 2023; Novack et al., 2024) that are related to chunked autoregression—to the best of our knowledge, these approaches have not been explored for live generation.

### v3

@@ 1 Introduction
Music exists in two complementary forms: as static recorded pieces (“music as a noun”), and as live performances collectively experienced in real time (“music as a verb”) [1]. This second form of live music is particularly tied to the fundamental human experiences of creative flow [2, 3], embodied expression [4], and social connection [5]. Despite this, modern generative AI systems for musical audio have had an overwhelming emphasis on offline, turn-based generation [6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
Open-weights models are particularly well-suited for live generative music because they can run locally on users’ devices. On-device inference has numerous benefits in music [17], enabling (1) lower latency by eliminating network requests, (2) higher reliability, facilitating usage in real-world contexts, (3) privacy guarantees, and (4) customization by artists. Cloud-based APIs address complementary use cases, offering access to models running on specialized hardware that are more powerful than those that would run on edge devices, at the cost of some of the benefits of open-weights models.
To allow users to navigate these tradeoffs based on their application goals, we introduce a pair of systems that span both paradigms: Magenta RT (open-weights, on-device) and Lyria RT (API, cloud-based). Both use the same core methodological framework, which centers around codec language modeling [18, 19] (Figure 1). Specifically, we train a language model (LM) to generate audio tokens from SpectroStream [20], using a method similar to [21, 22] to achieve live streaming [23].
We focus our subsequent discussion primarily on Magenta RT, which may be of higher interest to the AI research community for its ability to be finetuned with new controls [24] or explored for transfer learning [25].
Magenta RT offers first-of-its-kind live generation among open-weights models. On automatic metrics of music quality, it outperforms existing offline open-weights music generation models like MusicGen (Large) [9] and Stable Audio Open [14]. Moreover, Magenta RT uses 3838% fewer parameters (750750M) than Stable Audio Open (1.21.2B), and 7777% fewer than MusicGen Large (3.33.3B). Along with our codebase and model weights, we release a set of demos that run in real time on free-tier Colab TPUs (v2-8) and showcase three distinct use cases: live generation, finetuning, and a novel live audio input interaction that we call audio injection (Section 4.2). [^3] Audio samples and links to code, weights and demos are provided in the online supplement:
@@
Figure 1: Magenta RealTime is a live music model that generates an uninterrupted stream of music and responds continuously to user input. It generates audio in two-second chunks using a pipeline with three components: (1) MusicCoCa, a style embedding model, (2) SpectroStream [20], an audio codec model, and (3) an encoder-decoder language model. For each chunk, a style embedding is computed via a weighted average of MusicCoCa embeddings of text and audio prompts from the user. Given this style embedding and $10$ seconds ($5$ chunks) of past audio context, the language model decoder generates SpectroStream audio tokens for the new chunk, which is then decoded to audio.
@@ 2 Method
Magenta RT is a codec language model [18, 19] designed to generate high-fidelity stereo audio in real time based on acoustic style conditioning. To achieve this, we adopt a pipeline-based approach, using an encoder-decoder Transformer [26] to model audio tokens from SpectroStream [20] conditioned on style embeddings from our proposed MusicCoCa (Figure 1). See Appendix A for related work.
@@ 2.1 Audio Tokenization via SpectroStream
Here we adopt the recently proposed SpectroStream codec [20], a full-band ($f_{s}=48$kHz) multi-channel neural audio codec based on residual vector quantization (RVQ) [27], with an overall bandwidth of 1616kbps ($f_{k}=25$Hz, $d_{c}=64$, $|\mathbb{V}_{c}|=1024$). To facilitate live streaming, we reduce the bandwidth to 44kbps by generating only the first $16$ RVQ levels (coarse and medium from Figure 3), yielding a live throughput target of $400$ tokens per second. See Section C.1 for more details.
@@ 2.2 Style Embeddings via MusicCoCa
Our embedding model, MusicCoCa, builds upon MuLan [28] and CoCa [29]. It is a contrastive captioner (CoCa) consisting of two embedding towers, mapping each modality to a shared $768$-dimensional space. The audio embedding tower $M_{A}$ is a $12$-layer VisionTransformer (ViT) [30]. Its input is a log-mel spectrogram of a $10$s slice of 1616kHz audio ($128$ channels and length $992$; split into patches of size $16\times 16$). The text embedding tower $M_{T}$ is a $12$-layer Transformer, which operates on tokenized text with a maximum sequence length of $128$ tokens. We use attention pooling to reduce the activations of each tower to a single 768768d embedding, which can be subsequently quantized into $12$ discrete tokens with codebook size $|\mathbb{V}_{m}|=1024$. This tokenized representation (of audio or text) is then used as a conditioning signal in the LM encoder. Variable-length audio of duration $T$ may be embedded by zero padding to a multiple of $10$ seconds and then mean pooling across $\lceil\frac{T}{10}\rceil$ chunks.
In addition to the two embedding towers, MusicCoCa has a text decoder which can generate audio text captions. In our application this decoder is a shallow 3-layer Transformer which only serves a regularizing purpose. The MusicCoCa optimization objective, based on the CoCa framework [29], consists of contrastive and generative loss components which we weigh equally. We train the model using the Adafactor optimizer with a learning rate of $1\times 10^{-4}$ and 1,0001{,}000 warmup steps for a total of 16,00016{,}000 steps. The batch size is $1024$.
@@ 2.3 Modeling Framework
Magenta RT operates on audio tokens from a discrete audio codec, following an established practice in audio modeling [18, 31, 8, 9]. Our model autoregressively predicts discrete audio tokens conditioned on both preceding audio and a shared audio-text embedding. This modeling mechanism partly mirrors the MusicLM architecture [8], itself based on AudioLM [31]. Similarly to MusicLM, we use a Transformer-based architecture [26] for the token prediction model and enable text conditioning by leveraging a joint audio-text embedding model, using the target audio signal at training time, and text inputs at inference time. To facilitate live generation, we propose two high-level changes relative to MusicLM: (1) we replace the hierarchical cascade of multiple LMs with a single LM, using a recent method [23] similar to [21, 22] for efficiency, and (2) we propose a chunk-based autogressive approach to allow for infinite streaming generation.
@@ 2.3.1 Chunk-based autoregression with coarse context
In order to generate an infinite stream of audio, at inference time we must be able to predict an audio sequence with a length likely larger than what the model has seen during training. Such a mismatch in sequence length between training and inference is often found to result in unpredictable behavior and degraded performance [26, 32]. Previous work has addressed this issue via sliding attention windows with a relative positional encoding scheme [33, 34], informing the model about the relative distance between tokens instead of describing their absolute position within the sequence.
@@ 2.4 Encoder-Decoder Language Model
To model this distribution, we use an encoder-decoder Transformer [26] LM trained with T5X [35, 36]. We release pre-trained models using the T5  [35] Base and Large configurations.
@@
A key differentiating factor compared to prior work is our imposed constraint of achieving RTF $\geq 1\times$ generation to enable live interactive applications. Past work such as MusicLM [8] proposes a hierarchical cascade of language models to model tokens efficiently, while MusicGen [9] proposes a delay pattern—neither approach would achieve RTF $\geq 1\times$ for full bandwidth audio tokens. Instead, based on [23], our decoder comprises two connected Transformer modules. The “temporal” module constructs a temporal context by processing acoustic frames, where RVQ tokens within each frame are embedded and aggregated to yield a single frame-level embedding. The “depth” module then performs autoregressive prediction of the RVQ indices conditioned on the previous temporal context. With this setup, we achieve RTF=$1.8$ on H100 GPU with the T5 Large configuration.
@@ 3 Experiments
Live music models enable new types of interaction that are best assessed via direct human evaluation and extensive use. As such, many of our design choices were validated through play testing by team members and partner musicians, optimizing for how expressive and engaging the resulting instrument feels. A formalization of this interactive evaluation is presented in our user study in Appendix F.
We complement this subjective measure with more constrained experiments to examine the effects of our controls and provide comparisons to existing models where possible. We focus these experiments on assessing core capabilities that are common to both live and offline music audio generation models, such as audio quality and adherence to text conditioning (Section 3.2.1), alongside others that are unique to our live music models, such as the ability to generate musical transitions following changes in the conditioning signal (Section 3.2.2). In addition to our evaluation, we also note that, at the time of writing, Magenta RT is ranked as the top open-weights model on the Music Arena leaderboard [37], based on over 1k real-world user preferences.
@@ 3.1 Experimental Set-up
For inference, prompts (text or audio) are embedded and tokenized by MusicCoca using the first $6$ RVQ levels to match training. We sample with classifier-free guidance (CFG) [38, 39, 40], using a temperature of $1.3$, a top-K of $40$, and a CFG weight of $5.0$.
@@ 3.2 Results
Table 1: Instrumental music generation results on the Song Describer Dataset [41]. We compare to open models, using a fixed length of $47$s for all samples, though our models are capable of arbitrary-length generation. For all prior models, we report results from [14].
@@
Stable Audio Open [14] 44.1 kHz 1.1B 96.51 0.55 0.41
MusicGen-stereo-large [9] 32 kHz 3.3B 190.47 0.52 0.31
@@ 3.2.1 Audio quality and adherence to fixed text prompts
In this section, we compare Magenta RT to prior work under the offline text-to-music generation setting, where we keep the text conditioning fixed and sequentially generate chunks of audio up to a target length of $47$ seconds. While our model can generate audio of arbitrary length, we choose this duration for a fair comparison between models. Following the evaluation protocol in [14], we then assess the quality and text adherence of the resulting generations using three established metrics, the Fréchet Distance based on OpenL3 embeddings [42] (FDopenl3), the Kullback–Leibler divergence (KLpasst) and CLAPscore.
We show the results in Table 1. Magenta RT has the lowest FDopenl3 and KLpasst scores, indicating that the generated audio is plausible and closely matches the eval reference audio, including at the level of semantic correspondence [14]. The CLAPscore measures how well generated audio adheres to the specified text prompt. On this metric, Magenta RT scores between the other two models. The higher score of Stable Audio Open may be related to the fact that their model uses CLAP embeddings during training, as opposed to our model which trains using MusicCoCa.
@@ 3.2.2 Generating musical transitions
Live music models unlock the capability of responding dynamically to user inputs. To test this, we create a prompt transition evaluation. We pick pairs of text prompts and task the model with creating musical transitions between them. We linearly interpolate between MusicCoCa text embeddings of the start and end prompt over $60$ seconds ($6$ steps, $10$ seconds//step), and use this as style conditioning. We then measure cosine similarity between the audio embedding of the outputs and the conditioning embedding at that time. We perform transitions for $128$ prompt pairs (Appendix G).
@@ Appendix A Related Work
Our work on live music models connects to longstanding goals in computer music of live, computer-aided performance. Earlier work centered around score following: tracking live performances against known musical material to enable computer accompaniment [43, 44, 45], or systems that could both track input from human musicians and generate new symbolic material using simple probabilistic models [46, 47, 48, 49]. More recently, numerous systems have proposed live interaction with generative AI models of symbolic music [50, 51, 52, 53, 54], or small models trained on narrow music audio distributions [55, 56]. Here we aim to bridge the gap between the live interaction paradigm and the broad audio generation and control capabilities of large-scale generative AI models of music audio.
Many offline music audio generation models are based off of codec LMs [18, 19]. Codec LMs are theoretically capable of streaming provided they meet two criteria: (1) the language model and codec are causal (outputs never depend on future inputs), and (2) they generate with RTF $\geq 1\times$ for some chunk size $C$ (which lower bounds the control delay $D$). To the best of our knowledge, no existing codec LM music generation model [57, 6, 8, 9, 15, 58, 59] satisfies both criteria (most use non-causal codecs). Other music generation systems [7, 11, 12, 14, 13] are based off of latent diffusion—many achieve a throughput RTF $\geq 1\times$ but generate in a non-causal fashion. Some “outpainting” methods have been proposed for latent diffusion music models [60, 61] that are related to chunked autoregression—to the best of our knowledge, these approaches have not been explored for live generation.